OpenStudy (31356):

Factor Using Distributive Property (Double Check) -9^4y^3z^5-6x^4yz^6+3x^4z^6 MEDAL REWARDED!!

OpenStudy (31356):

My guess is 3x^4z^5(-3y^3-2zy+z)

OpenStudy (31356):

@Mertsj

OpenStudy (31356):

@jim_thompson5910

OpenStudy (mertsj):

Good job!!

OpenStudy (31356):

So it's correct?

OpenStudy (31356):

Thanks for the medal @jigglypuff314

jigglypuff314 (jigglypuff314):

^_^ only thing that threw me off was you were missing an x in your question XD 9^4 ?!?!

OpenStudy (31356):

Oh yea forgot to include that :)

OpenStudy (31356):

Can both of you help me double check this?

jigglypuff314 (jigglypuff314):

of course :)

OpenStudy (31356):

35-28n+21 = -28(n-2) right?

jigglypuff314 (jigglypuff314):

yep! :)

OpenStudy (31356):

Yea! Can you double check 4 more please? :)

jigglypuff314 (jigglypuff314):

lol k XD

OpenStudy (31356):

1) 10x^6+8x^3= 2x^3(5^3+4) 2)5+5n= 5(n+1) 3)-25p+5= -5(5p-1) 4)12n^2-9n= 3n(4n-3)

OpenStudy (31356):

Thanks in advance! :D

jigglypuff314 (jigglypuff314):

you forgot to type an x in the answer for the first one :P other than that Great Job!!! :D

OpenStudy (31356):

Oh sorry again, and thanks for double checking! So far that's all! :D

OpenStudy (31356):

Have you posted other flute music on SoundCloud?

jigglypuff314 (jigglypuff314):

yup! ^.^ I'll post more tomorrow too (working on an Etude Collection :) ) and I have a thing open in music right now :P http://openstudy.com/study#/updates/52fbf19fe4b03b1616efcc1f

OpenStudy (31356):

Cool!! :P